40 Use the method of completing the square to transform the quadratic equation into the equation form (x -p2-q. ?9 + 12x-6x2 =0 A) (x-1)2??0.5 B) 12- -1.5 C) D) (1)2-1.5 (x-1)2-0.5
Answer #1

Given equation is-
\small -9+12x-6x^2=0

Divide both sides by \small -6
\small =>\frac{-9+12x-6x^2}{-6}=0
\small =>\frac{-9}{-6}+\frac{12x}{-6}-\frac{6x^2}{-6}=0
\small =>\frac{3}{2}-2x+x^2=0

rearanging-
\small =>x^2-2x+\frac{3}{2}=0

Split the term \tiny \frac{3}{2}=\frac{2+1}{2}
\small =>x^2-2x+\frac{2+1}{2}=0
\small =>x^2-2x+\frac{2}{2}+\frac{1}{2}=0
\small =>x^2-2x+1+\frac{1}{2}=0

As we know \small (a-b)^2=a^2-2ab+b^2
\small =>x^2-2x+1^2+\frac{1}{2}=0 (as we know 12=1)
\small =>(x-1)^2+\frac{1}{2}=0

Subtract both sides by \tiny \frac{1}{2}
\small =>(x-1)^2+\frac{1}{2}-\frac{1}{2}=-\frac{1}{2}
\small =>(x-1)^2=-\frac{1}{2}
\small =>(x-1)^2=-0.5

Hence answer is-
\small A)\quad (x-1)^2=-0.5